OpenStudy (puppy730):

William is putting up signs on the edge of the road every of a mile. He just put a sign down and he has of a mile of road left in which to place his signs. He wants to determine how many more signs he should put on the edge of the road. William should put up more signs.

OpenStudy (danjs):

every..what fraction of a mile?

OpenStudy (puppy730):

William is putting up signs on the edge of the road every 1/6 of a mile. He just put a sign down and he has 7/8 of a mile of road left in which to place his signs. He wants to determine how many more signs he should put on the edge of the road. William should put up more signs.

OpenStudy (danjs):

ok he has 7/8 mile to go and places a sign every 1/6 mile

OpenStudy (danjs):

\[\frac{ \frac{ 7 }{ 8 } }{ \frac{ 1 }{ 6 } }\]

OpenStudy (puppy730):

yes

OpenStudy (danjs):

\[\frac{ 7 }{ 8 } * \frac{ 6 }{ 1 }\]

OpenStudy (danjs):

\[\frac{ 42 }{ 8 } \approx 5.25\]

OpenStudy (danjs):

so he has enough road left for 5 signs

OpenStudy (puppy730):

ok ty very much